OpenStudy (anonymous):

A publicist's compensation package includes the total cost of a $180-per-month health insurance plan, the total cost of a $35-per-month life insurance plan, and a salary of $42,000 per year. What is the yearly value of the compensation package?

OpenStudy (texaschic101):

180 per month (for 1 yr) = 180 * 12 = 2160 ( I am multiplying by 12 because there is 12 months per year. 35 per month (for 1 yr) = 35 * 12 = 420 salary (per yr) = 42,000 total per yr = 2160 + 420 + 42,000 = 44,580 I think....I am assuming the compensation package includes the salary